I have a given volume of 300cm³. The ratio between height and radius is 1.1:1 but I can't calculate the inverse formula for the volume. Please help
Answer by venugopalramana(3286)

I have a given volume of 300cm³. The ratio between height and radius is 1.1:1 but I can't calculate the inverse formula for the volume. Please help
WHAT IS IT ? CYLINDER OR CONE?
LET HEIGHT =1.1X...AND RADIUS =1X..AS THEIR RATIO IS 1.1:1
IF CONE ...V=PI*R*R*H/3
300=PI*X*X*1.1*X/3
X^3=3*300/(1.1*PI)
X= 6.387
IF CYLINDER....V=PI*R*R*H...YOU MAY TRY YOUR SELF